(1 point) Test the claim that the population of sophomore college students has a mean grade point average of 2.24 versus a meDETAILED WORK PLEASE
Answer #1

Solution :

Given that,

Population mean = \mu = 2.24

Sample mean = \bar x = 2.38

Sample standard deviation = s = 0.5

Sample size = n = 15

Level of significance = \alpha = 0.01

This a right (One) tailed test.

The null and alternative hypothesis is,  

Ho: \mu \leq 2.24

Ha: \mu > 2.24

The test statistics,

t = ( \bar x - \mu )/ (s/\sqrt{n})

= ( 2.38 - 2.24 ) / ( 0.5 / \sqrt{} 15 )

= 1.084

Critical value of  the significance level is α = 0.01, and the critical value for a right-tailed test is

t_{c} = 2.624

Since it is observed that t = 1.084 \leq t_{c} = 2.624, it is then concluded that the null hypothesis is fail to rejected.

Conclusion:

A)